Craig is 24 years younger than his father . In 10 years , Craig's father will be three timesas old as Craig will be. Find their ages now.
Craig=c-24 c+24=d 3(c+10)=d+10
Dad= d 3c+30=d+10
3c+20=d

Answer by darq(90)

Craig is 24 years younger than his father . In 10 years , Craig's father will be three timesas old as Craig will be. Find their ages now.

c = f - 24
3*(c + 10) = f + 10

3c = 3f - 72
3c + 30 = f + 10

3c = 3f - 72
3c = f - 20
f - 20 = 3f - 72
2f = 52
father = 26
craig = 26 - 2
craig = 2